OpenStudy (anonymous):

8.) If a triangle has sides of lengths 9, 12, and 15, is it a right triangle? Yes No Sometimes Not enough information

OpenStudy (dumbcow):

if 9^2+12^2 = 15^2 then yes otherwise no

OpenStudy (anonymous):

12^2+ 9^2=15^2 225=225 yes